OpenStudy (anonymous):

write the equation for a line in standard form that goes through the points -1,3 and -2,5

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

y-y1/x-x1= y2-y1/x2-x1 y-3/x+1 =2/-1 y-3 =-2x -2 y+2x =1 thats the required eqn

OpenStudy (anonymous):

?

OpenStudy (anonymous):

wht u didnt get it?

OpenStudy (anonymous):

it's y = -2x +1

OpenStudy (anonymous):

i though it was -2x but you have 2x

OpenStudy (anonymous):

but in my case dear the 2x is on the other side of the = sign

OpenStudy (anonymous):

so your right and im wrong?

OpenStudy (anonymous):

yea stom i was typing and didn't see you had responded hahaha sorry

OpenStudy (anonymous):

oh u dont have to be sorry , :) btw @devi have u got the working?